OpenStudy (kevon18):

Find all of the zeros and write a linear factorization of the function. f(x)=x^4+x^3+4x^2-4x-32

OpenStudy (3mar):

May I help?

OpenStudy (kevon18):

Yes

OpenStudy (tgstudios):

this?\[f(x)=x ^{4}+x ^{3}+4x ^{2}-4x-32\]

OpenStudy (kevon18):

f(x)=x^4+x^3+4x^2-4x-32

OpenStudy (3mar):

@TGstudios it is your turn.

OpenStudy (tgstudios):

XD... this is calulus1LOL... I have no clue what calculus is!!

OpenStudy (3mar):

No need to calculus! It is just factorizing by grouping, no more!

OpenStudy (kevon18):

This is Pre-Calculus Work!

OpenStudy (3mar):

Can you proceed?

OpenStudy (sshayer):

\[factors~ of ~32 ~are~ \pm1,\pm2,\pm4,\pm8,\pm 16 \] try x=2, \[f(2)=2^4+2^3+4*2^2-4*2-32=16+8+16-8-32=0\] so x=0 is a zero of f(x)